[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Sep 21, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Sep 21 19:37:56 UTC 2016


HSU Shift Log for Wed Sep 21 2016 15:37:50 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: TSB requested that we not reboot any NAWIPS boxes  
tonight since there may be an issue with the video driver due to the latest  
patch that was installed. If you have NAWIPS workstation issues tonight,  
please call the SDM.

Other Techincal issues: Text TWOs did not update on the webpage at 18Z.  
GTWO ran fine. Dave Z. said there were issues with dataflow caused by  
issues outside of NHC. See his email for more details.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160921/012badc1/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list